I need help w/step by step instruction on how to solve this word problem.
The speed of a boat in still water is 10 km/h. The boat travels 48km upstream and 48km downstream in a total time of 10 hr. What is the speed of the stream?
I think I know the answer but am unsure of how to show it, step by step, in an algebraic equation.
Thanks!

Answer by Alan3354(69443)

The speed of a boat in still water is 10 km/h. The boat travels 48km upstream and 48km downstream in a total time of 10 hr. What is the speed of the stream?
----------------
Use the d = rt, distance = rate x time.
d = rt
t = d/r
r = rate of the boat
c = speed of the current
----------------
Upstream, t = 48/(r-c) current subtracted from boat speed
Downstream, t = 48/(r+c)
48/(10-c) + 48/(10+c) = 10
Now it's just another equation.
48(10+c) + 48(10-c) = 10*(10+c)*(10-c)
960 = 1000 - 10c^2
c^2 = 4
c = 2 km/hr